The following fractions represent just three different numbers. Separate them into three groups of equivalent fractions, by changing each one to its simplest form.
12/72

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

12 / 72 = (1 × 12) / 6 × 12)
=1 / 6

​Thanks